40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5542     Entry time: Mon Sep 26 11:35:44 2011
Author: steve 
Type: Update 
Category: Cameras 
Subject: arms cameras upgraded 

The arm's CCD cameras were reset as picture shows last week.

The height adjustment only works at ITMX. Short post holders are ordered to make this feature  available on the rest.

The 75 ohms video and power supply cables are stress relieved. Solid cover can be attached now without  miss aligning cameras.

Attachment 1: P1080251.JPG  2.035 MB  | Hide | Hide all
P1080251.JPG
ELOG V3.1.3-